Reports 1-1 of 1 Clear search Modify search
AdV-SGD (FDS commissioning)
delauren, grimaldi, sorrentino, vardaro - 13:02 Saturday 23 October 2021 (53560) Print this report
Improvement of ModeMatching between GR and FC

The aim of this shift was to increase the Mode Matching  between the Green beam and the filter cavity. We worked only with EQB1_GLAOM and we reduced the Mode Mismatch to less than 8.0 %  (gps 1318946495 14:01:17 UTC). [Tem00 7.85; LG10 0.7 V]

We mounted the EQB1_GAOM  on a motorized slit in order to have a fine control of its position. After that we re-aligned the AOM in double pass configuration. 

We did not connect the slits to the DAQ system, but we tuned the position by hand, optimizing the resonance peak of the TEM00.  Before checking the high of the 2nd order peak we optimized the alignment using the Drift control and the BPC.

At the end off activities in DET lab, we lost the alignment of the Green beam: the power drop from ~7.6V to 4V. 

In order to recover the GR alignment we had to remove the set point of the Drift control, so the GR Beam now is lower on FCEB_GR_CAM (Figure 3). During the alignment activities SQB1 unlocked and it was recovered by the operator.

At the end of the Recovery we realigned the IR beam and we check the MM of the GR beam: we measured ~8% of Mode Mismatch (Figure 1 and Figure 2)

Images attached to this report
Search Help